OpenStudy (anonymous):

Will had $15 and fran had $45. Will spent $3 at the store. Fran spent an equivalent fraction of her money. How much did fran spend?

OpenStudy (anonymous):

You can set up their money like this: \[w = 15\]\[f = 45\] Then, you can simplify them like this: \[w = \frac{ 3 }{ 15 } = \frac{ 1 }{ 5 }\] Now you know that if fran spent an equivalent fraction, the simplified versions should be the same: \[w = \frac{ 1 }{ 5 } = f = \frac{ x }{ 45 }\]\[\frac{ 1 }{ 5 } = \frac{ x }{ 45 }\] Then you can simplify by multiplying both things by the bottom of the fraction with x in it! \[\frac{ 45 }{ 1 }*\frac{ 1 }{ 5 } = \frac{ 45 }{ 5 } = 9\] Understand?

OpenStudy (anonymous):

Thank you I thought the same but wasn't sure

OpenStudy (anonymous):

No problem.